The diameter of an approximately spherical May Day ball is 35 cm. The balloon is filled with helium so that the balloon has an overpressure of 0.12 bar at a temperature of 18°C. What is the mass of helium in the sphere?
1
Expert's answer
2020-05-05T18:41:25-0400
"pV=\frac{m}{M}RT\to \frac{4}{3}\pi r^3p=\frac{m}{M}RT"

"\frac{4}{3}\pi (0.35)^3(0.12\cdot 10^5)=\frac{m}{4}(8.31)(273+18)"

"m=3.6\ g"